Apache Point Observatory
New Mexico, USA
ARC 3.5-m Instrument history
Past problems and technical issues which our users should be aware of when reducing archival data. This page is in construction. More dates will be added for each instrument as time allows. If you have specific questions regarding your data and the status of the instrument in the past contact APO for more information.
AGILE
ARCES
- ARCES Users Guide Appendix B. Instrument History
- 2012-09-05: Reports of varying levels in the CuS (blue) Flats with no change in the unfiltered Qth Flats caused me to suspect the bulb is starting to fail. Changed out the bulb in the QTH lamp today.
- 2012-08-10: The new MR152 grating was installed and aligned over summer shutdown. Please be aware not to use old cals or wavelength solutions since the wavelengths in each order have shifted. Contact the Observing Specialists or the Telescope Engineer for more details.
- 2012-06-22: Replaced Red and Blue slitviewer filters with new ND4 and ND5 filters. User guide updated with more information on the ND filters.
- 2012-03-04: Echelle slitviewer masks fixed
DIS
- DIS Users Guide Appendix B. Instrument History
- 2012-08-28: Yesterday we added an Aluminum Nitrade electrical isolation disk between the cold head and cold strap block inside the dewar. The Red camera was pumped on for 24 hours and cooled this morning. Initial tests show that the biases look normal again and that dark current is low (< 0.02 DN/sec).
- 2012-08-24: The Red camera is experiencing excessive noise as in readout, possibly due to extra dark current. It was taken offline today and warmed in preparation for a more thorough investigation next week.
- 2012-08-10: Over summer shutdown both DIS cameras and cryotigers were fully serviced. The Blue camera with a new cryogenic cold head. Both camera showed water vapor spots on their field lenses. They were removed and cleaned. There was minor damage done to the Red camera field lens when it was reinstalled.
- 2012-05-07: DIS Blue contamination shown to be gone.
- 2012-05-04: DIS Blue retruned to service late in the day. Unable to determine yet is contamination is still present.
- 2012-04-30: DIS removed from service today to deal with contamination on the blue camera. Expect it to be out of service until Thursday of this week
- 2012-04-24: DIS blue is tracking ambient temperatures and is running about 7 - 10 degrees warm (-87C). Extended wings on the spectrum suggest a scattering source on optics or the ccd. Will monitor and address if it continues to worsen. Otherwise, servicing this summer is planned.
- 2012-03-28: NEW B400 grating installed and aligned today.
- 2012-03-26: DIS Blue camera ion pump replaced. It is warm and still needs more pumping this week. It is not in service until further notice. DIS B400 Grating has been removed for swap with new grating.
- 2012-02-17: DIS Blue camera temperature is back to normal.
- 2012-02-16: DIS Blue camera vacuum has gone soft and is experiencing temperature regulation issues today. We are putting it on the vacuum pump. THis will take DIS out of service until the vacuum issue is resolved. If all goes well it will still b e available for this evening.
- 2012-02-10: DIS Users Guide updated today with new recommended exposure times for flat fielding using the truss lamps.
- 2012-02-09: New recommended exposure times for flat field lamps can be found here.The DIS User Guide will be updated shortly.
NICFPS
- NICFPS Users Guide Appendix B. Instrument History
- 2012-08-10: The NICFPS ICS Hard drive failed bringing the instrument back up. Fritz was able to install a new hard drive and reinstall the most current version of the software. A back-up was done at the same time.
SPICAM
- SPICAM Users Guide Appendix B. Instrument History
- 2012-09-28: Spicam is back with its new ion pump and controller. There is an electrical noise problem seen in biases and darks. Looks like a grounding problem from the ion pump through the coldhead and cold strap. Will address asap.
- 2012-09-21: The ion pump for SPIcam failed this evening. The instrument is warming for a replacement. It will not be available for use until this repair is done
- 2012-08-28: The cryotiger was serviced over shutdown. Temperatures continue to be running consistently 10 - 15 degrees warmer than expected. APO has a call in with Brooks Polycold for input.
- 2012-05-17: Returned to service today. Still needs an instrument block due to the amount of disassembly that was performed.
- 2012-04-26: Removed from service for maintenance
- 2012-04-24: Servicing scheduled starting Thursday April 26. The instrument will be unavailable after that date until further notice.
- 2012-04-09: SPicam's cold head appears to be partially restricted, limiting its cooling capacity. It is running 20 degrees above nominal. We therefore strongly recommend that users take darks to accompany their data. The following page gives more information of the problem. Plans are in progress to replace the coldhead the end of April and beginning of May.
- 2012-04-09: SPicam's cold head appears to be partially restricted, limiting its cooling capacity. It is running 20 degrees above nominal. We therefore strongly recommend that users take darks to accompany their data. The following page gives more information of the problem. Plans are in progress to replace the coldhead the end of April and beginning of May.
- 2012-04-03: Instrument returned to service.
- 2012-03-30: Instrument removed from service while we service the Cryotiger, lines and coldhead
- 2012-03-28: Camera partially warmed again. Still usable but darks are recommended. Will address the cryogenics system in the next few days
- 2012-03-06: Camera again cold. Contamination appears to be gone.
- 2012-03-01: Warmed and placed on vacuum pump to remove contamination. Not available until next week when we can take off pump and cool.
- 2012-02-27: Contamination on the CCD has reappeared. We were able to get the frost to partially sublimate off the chip by raising the chip heater a few degrees. Both visual and flat fields show evidence of contamination. About 400 counts above background. We are attempting to reduce the amount further however we will need several days to warm and pump on the instrument. For an example of what we are seeing look here.
- 2012-02-09: Spicam User Guide updated with new recommended exposure times for flat fielding using the truss lamps.
- 2012-01-24: SPIcam exhibits a low level pattern noise as seen in biases. This noise is very intermittent and has existed for the life of this instrument. Cause has never been determined. Examples of the SPIcam pickup noise as seen in biases can be seen here.
TRIPLESPEC
- TRIPLESPEC Users Guide Instrument information
- 2012-11-01: New instrument cart is now in full use. It was causing interference with the lip on the floor below the NA2 rotator so we cut a section of that kickplate lip away.
- 2012-10-20: Temperature logging fixed but not reported to APO however in the meantime critical alerts were off. They are back on now.
- 2012-09-21: Found the instrument was not properly logging temperatures. Power cycled the aux box for Matt Nelson. It is back up
- 2012-08-10: A reminder to users. Since the slitwheel is locked in place, True instrument darks are not possible at this time.
- 2012-02-19: The slitwheel mechanism has been powered off and disabled at this time. We are working with the UVa Instrument team to develop a repair plan. It currently is locked in the 1.1 arcsecond slit position.
- 2012-02-10: Exposure times for flat fielding using the truss lamps have not changed with the new lamps for Triplespec.
- 2012-02-10: Triplespec slit wheel has been experiencing problems where it freezes position or is sluggish to respond to commands. We are afraid that it will move to a non slit position and render the instrument unusable. As a result the engineers at UVa and APO staff are recommending the slit not be moved from the default 1.1 arcsecond slit, at least until we have a better handle on what is going on.
- 2012-02-07: Tripespec was exhibiting a low level intermittent pattern noise that can be seen in FS=8. It is on the order of +/- 20 DN noise. It has also been seen at lower levels (factor of 4) in FS=1 going back to April, 2010. This noise has been identified as getting worse than a few DN peak to peak as recently as this past December. For more information look here. APO is actively working on resolving this problem. We believe we may have made improvements today but it may take afew days to a week of monitoring to tell. Contact us if you see this in data after this date.
Misc